# MS-Occ: Multi-Stage LiDAR-Camera Fusion for 3D Semantic Occupancy Prediction

### v1

@@ Abstract
Accurate 3D semantic occupancy perception is essential for autonomous driving in complex environments with diverse and irregular objects. While vision-centric methods suffer from geometric inaccuracies, LiDAR-based approaches often lack rich semantic information. To address these limitations, MS-Occ, a novel multi-stage LiDAR-camera fusion framework which includes middle-stage fusion and late-stage fusion, is proposed, integrating LiDAR’s geometric fidelity with camera-based semantic richness via hierarchical cross-modal fusion. The framework introduces innovations at two critical stages: (1) In the middle-stage feature fusion, the Gaussian-Geo module leverages Gaussian kernel rendering on sparse LiDAR depth maps to enhance 2D image features with dense geometric priors, and the Semantic-Aware module enriches LiDAR voxels with semantic context via deformable cross-attention; (2) In the late-stage voxel fusion, the Adaptive Fusion (AF) module dynamically balances voxel features across modalities, while the High Classification Confidence Voxel Fusion (HCCVF) module resolves semantic inconsistencies using self-attention-based refinement. Experiments on the nuScenes-OpenOccupancy benchmark show that MS-Occ achieves an Intersection over Union (IoU) of $32.1\%$ and a mean IoU (mIoU) of $25.3\%$, surpassing the state-of-the-art by $+0.7\%$ IoU and $+2.4\%$ mIoU. Ablation studies further validate the contribution of each module, with substantial improvements in small-object perception, demonstrating the practical value of MS-Occ for safety-critical autonomous driving scenarios.
Index Terms:
3D semantic occupancy, LiDAR-camera fusion, multi-stage fusion, autonomous driving, sensor fusion, deformable attention.
@@ I Introduction
3D semantic occupancy plays a vital role in autonomous driving by enabling a comprehensive understanding of complex environments, which is essential for downstream tasks such as vehicle planning and control. In contrast to traditional 3D object detection, occupancy perception captures the spatial distribution of both irregular and non-rigid objects, extending beyond predefined object categories. Vision-based 3D semantic occupancy perception has recently demonstrated notable performance across various technical paradigms, including monocular setups [1], surround-view camera systems [2, 3, 4], and efficient architectural designs [5, 6]. However, vision-based methods encounter inherent challenges, including inaccurate depth estimation, limited geometric fidelity, and sensitivity to lighting conditions, which collectively undermine robustness in real-world driving environments. These challenges highlight the importance of integrating LiDAR and camera data to improve occupancy perception. Although 4D radar has been explored in recent studies [7, 8] as a supplement to mitigate some of these issues, it remains constrained by limitations such as point cloud noise and sparsity.
@@
Figure 1: Comparison of occupancy fusion stages in encoder design. (a) Middle-stage fusion (e.g., [9]), where 2D image features are directly fused with 3D LiDAR voxel features. (b) Late-stage fusion, including methods such as [10, 11, 12, 13], which explicitly lift 2D image features to 3D voxel features (e.g., using [14]), followed by 3D fusion in voxel space. (c) Multi-stage fusion, as proposed in our MS-Occ, which integrates both middle and late-stage fusion within a hybrid architecture.
@@
The integration of LiDAR and camera technologies has attracted increasing research attention in the domain of multi-modal sensor fusion. Building on insights from [15, 16], multi-modal 3D semantic occupancy perception can be classified into four fusion stages based on the timing of feature integration: early-stage fusion, middle-stage fusion, late-stage fusion, and multi-stage fusion. Early-stage fusion involves the combination of raw or preprocessed sensor data, allowing for full utilization of the original information. However, it is prone to misalignment caused by modality-specific limitations and calibration errors. As a result, most existing methods prioritize middle or late-stage fusion, as illustrated in Fig. 1(a) and (b).
@@
Middle-stage fusion integrates 2D image features with 3D LiDAR features through projection-based fusion $E_{\text{fusion}}^{\text{3D}\rightarrow\text{2D}}$ or cross-attention mechanisms $E_{\text{fusion}}^{\text{2D}\rightarrow\text{3D}}$. For instance, OccLoff [9] employs cross-attention to fuse 2D image features with LiDAR voxel features. Nonetheless, middle-stage fusion methods often fail to fully exploit the strengths of each modality within the 3D voxel space. In contrast, late-stage fusion combines 3D image voxel features, explicitly lifted from 2D image features ($E_{\text{imgs}}^{\text{3DLift}}$), with 3D LiDAR voxel features in voxel space ($E_{\text{fusion}}^{\text{3D}}$). Representative examples include M-CONet [10] and Co-Occ [11], both of which enhance performance through late-stage fusion. OpenOccupancy [10] presents a novel LiDAR-camera fusion baseline that leverages the complementary advantages of the sensors and establishes the first benchmark for fusion-based surrounding occupancy perception. Co-Occ [11] further refines LiDAR features by incorporating neighboring camera features using a kNN search strategy in GSFusion.
@@
Despite their success, these approaches perform fusion solely in voxel space, overlooking the hierarchical nature of 2D image features. This omission introduces geometric and semantic biases during voxel grid construction, thereby limiting the alignment of multi-modal features in the 3D voxel domain. To address these limitations, multi-stage fusion, illustrated in Fig. 1(c), combines both middle- and late-stage fusion strategies. This method leverages 2D visual and 3D LiDAR spatial features to achieve robust alignment of semantic richness and geometric precision, thereby further enhancing the final 3D semantic occupancy perception performance. Nevertheless, multi-stage fusion remains underexplored in existing research.
@@
This study introduces MS-Occ, which, to the best of the authors’ knowledge, is the first LiDAR-camera fusion framework to perform 3D semantic occupancy prediction using multi-stage fusion. Initially, features are extracted from both LiDAR point clouds and surround-view images. These features are then enhanced through middle-stage feature fusion, where 2D camera features are enriched with geometric information using the proposed Gaussian-Geo module, and LiDAR features are augmented with semantic cues via the Semantic-Aware module. This fusion stage utilizes two novel modules, leveraging rich intermediate representations to address the inherent limitations of individual sensing modalities. In the Gaussian-Geo module, a Gaussian kernel is applied to produce a denser LiDAR heatmap, which interacts with the original 2D image features to generate geometrically enhanced camera features. These enhanced features are subsequently fused with 3D LiDAR features using deformable attention in the semantic-aware module, thereby injecting semantic context into the LiDAR representations.
@@
Following this middle-stage fusion, a Late-stage voxel fusion is performed to further refine the fused voxel features from both modalities. This stage comprises two complementary components: the High Classification Confidence Voxel Fusion (HCCVF) module and the Adaptive Fusion (AF) module. The AF module initially performs adaptive fusion of voxel features, following the fusion method of M-CONet [10], to integrate voxel representations from both LiDAR and camera streams.
The innovative HCCVF Module then selectively refines high-confidence features from each modality through an attention mechanism, thereby mitigating semantic ambiguities between them. The final fused voxel features are passed through a prediction head to generate the 3D semantic occupancy output.
@@
  • A novel LiDAR-camera multi-stage fusion framework for 3D semantic occupancy prediction is proposed. To the best of the authors’ knowledge, this is the first multi-stage fusion architecture designed for the 3D semantic occupancy prediction task. The proposed framework efficiently processes LiDAR and camera data to generate 3D voxel representations that compensate for modality-specific limitations, thereby improving performance.
  • Four dedicated modules are introduced to facilitate unified and fine-grained 3D semantic occupancy perception. The Gaussian-Geo module, employed during the middle-stage fusion, enriches camera features with dense geometric priors from LiDAR using Gaussian rendering. The Semantic-Aware module transfers image semantics to LiDAR voxels through deformable attention. In the late-stage fusion, the HCCVF and AF modules are integrated: HCCVF resolves semantic ambiguities via self-attention, while the AF module dynamically balances geometric and semantic voxel features.
  • Extensive experiments on the nuScenes-OpenOccupancy benchmark demonstrate that MS-Occ outperforms existing state-of-the-art methods, achieving superior performance in both geometric and semantic accuracy. Notably, the framework achieves 32.1% IoU and 25.3% mIoU, with significant improvements in safety-critical categories such as pedestrians, bicycles, and small objects.
@@ II-A Vision-based 3D Occupancy Prediction
In the domain of vision-based occupancy perception, significant advances have been made in recent years. The transformation of 2D image features into 3D representations to generate Bird’s-Eye View (BEV) features [17], Tri-Perspective View (TPV) features [18], and voxel features [2] is critical for vision-based 3D occupancy prediction. Three dominant approaches have emerged for this transformation [19]: (1) The projection-based method [1, 20] employs perspective projection models to establish geometric correspondences by projecting voxel centroids from 3D space onto 2D image planes, thereby enhancing the geometric awareness of 2D image features. (2) The depth prediction method [21, 17, 6], derived from the Lift-Splat-Shoot (LSS) paradigm, uses estimated depth distributions to lift 2D features into 3D space. (3) The cross-attention method [2, 22, 18] facilitates feature interaction between 3D voxel grids and 2D feature maps through deformable attention [23], enabling adaptive spatial feature alignment. Recent results have improved computational efficiency by employing lightweight models, such as SparseOcc [5] and EFFOcc [13].
However, vision-based 3D occupancy perception faces fundamental limitations due to sensitivity to illumination changes, inherent depth estimation errors, and compromises in geometric precision and efficiency.
@@ II-B LiDAR-Camera Fusion for 3D Occupancy Prediction
Recent advancements in LiDAR-camera fusion have demonstrated exceptional performance across 3D occupancy perception tasks through four primary strategies [19]: (1) Concatenation-based methods [24, 11], which concatenate 3D feature volumes from distinct modalities along either feature or sequence dimensions, followed by convolutional layers to enable cross-modal feature interaction. (2) Summation-based approaches [10, 12], which perform element-wise additive operations to combine geometrically aligned sensor representations using learnable modality-specific weighting coefficients. (3) Attention-based methods [9], inspired by SparseFusion [25] and TransFusion [26], which apply cross-modality attention mechanisms [23] to establish latent correspondences between LiDAR features and image representations. (4) Projection-based methods [27, 28], which align LiDAR points with image pixels via calibrated projection to facilitate 2D feature fusion.
However, the aforementioned strategies primarily focus on either middle- or late-stage fusion mechanisms, often neglecting the importance of fusion timing in optimizing the integration of modality-specific representations. This oversight constrains the potential for dynamic alignment of representations across modalities during optimal integration phases. In this letter, the proposed MS-Occ addresses this limitation by leveraging modality-specific features at multiple fusion stages to facilitate cross-modal alignment and enhance the quality of fused voxel-level representations.
@@ III Method
The proposed MS-Occ leverages the complementary strengths of LiDAR and camera modalities to enhance 3D semantic occupancy prediction. The overall architecture is illustrated in Fig. 2. Modality-specific features are acquired using 2D and 3D encoders.
@@ III-A1 Gaussian-Geo Module
Figure 3: Illustration of intermediate outputs from the proposed Gaussian-Geo module. The raw LiDAR point cloud is densified using Gaussian rendering, and the resulting geometric information is subsequently transferred to the camera modality to enhance its spatial representation.
@@
The Gaussian-Geo module is designed to incorporate geometric priors through Gaussian rendering, drawing inspiration from recent approaches [29, 30] that utilize 3D Gaussians [31] for scene representation, as illustrated in Fig. 3. Specifically, depth maps are initially obtained from LiDAR point clouds. A simple approach involves directly projecting point clouds onto images using intrinsic and extrinsic calibration parameters to generate depth maps. However, due to the inherent sparsity of LiDAR data, the resulting depth maps are typically incomplete and lack detailed features. To address this issue, a Gaussian kernel rendering technique is introduced to enhance the projected depth maps across multi-view images, resulting in improved depth representations $Y\in[0,D_{\text{pt}}]^{H_{\mathrm{C}}\times W_{\mathrm{C}}\times 1}$, where $D_{\text{pt}}$ denotes the LiDAR point depth value, $H_{\mathrm{C}}$ and $W_{\mathrm{C}}$ represent the height and width of the original camera image, respectively. The raw LiDAR points are splatted onto multi-view images. For each hit position $p_{j}\in\mathbb{R}^{\mathrm{2}}$ as a center, a Gaussian kernel with the corresponding depth value is generated in the image plane. The final depth value at position $q\in\mathbb{R}^{\mathrm{2}}$ is computed as:
@@
$Y_{q}=\underset{j\in\mathcal{G}_{\text{hit}}}{\min}\left(D_{\text{pt}}^{j}\mathtt{exp}\left(-\frac{\left(q-p_{j}\right)^{2}}{\mathrm{2\sigma^{2}}}\right)\right),$ (1)
@@
where $\mathrm{\sigma}$ controls the spread of the kernel, $\mathcal{G}_{\text{hit}}$ represents the set of overlapping Gaussian kernels at position $q$, and $j$ indexes the overlapping kernels.
@@
In regions with multiple overlapping kernels, the minimum value is selected to generate a denser and more continuous multi-view depth map. The enhanced depth maps $Y$ are then passed through a shared encoder $\mathrm{\varphi\left(\cdot\right)}$ to extract depth features $F_{\mathrm{D}}^{(m,i)}\in\mathbb{R}^{H_{\mathrm{I}}\times W_{\mathrm{I}}\times C_{\mathrm{D}}}$, where $C_{\mathrm{D}}$ denotes the dimension of the depth features. These features are concatenated with the corresponding 2D image features $F_{\mathrm{I}}^{(m,i)}$, and the composite features are processed by a depth fusion network $\mathrm{\psi\left(\cdot\right)}$ to obtain depth-aware 2D image features $F_{\mathrm{D2I}}^{(m,i)}\in\mathbb{R}^{H_{\mathrm{I}}\times W_{\mathrm{I}}\times C_{\mathrm{I}}}$, thereby enhancing the geometric understanding of the image content. The overall process can be formulated as:
$F_{\mathrm{D2I}}^{(m,i)}=\mathrm{\psi}\left(\mathtt{Concat}\left[F_{\mathrm{D}}^{(m,i)},\mathit{F_{\mathrm{I}}^{(m,i)}}\right]\right),$ (2)
where
$F_{\mathrm{D}}^{(m,i)}=\mathrm{\varphi}\left(Y\right).$ (3)
@@
Ultimately, the depth-aware 2D image features are fed into the LSS [14] module, which transforms them into 3D image voxels $V_{\mathrm{I}}\in\mathbb{R}^{D\times H\times W\times C}$, matching the dimensions of the LiDAR voxel features. This transformation enriches 2D image features with geometric context, yielding a more comprehensive and spatially accurate scene representation.

### v2

@@ Abstract
Accurate 3D semantic occupancy perception is essential for autonomous driving in complex environments with diverse and irregular objects. While vision-centric methods suffer from geometric inaccuracies, LiDAR-based approaches often lack rich semantic information. To address these limitations, MS-Occ, a novel multi-stage LiDAR-camera fusion framework which includes middle-stage fusion and late-stage fusion, is proposed, integrating LiDAR’s geometric fidelity with camera-based semantic richness via hierarchical cross-modal fusion. The framework introduces innovations at two critical stages: (1) In the middle-stage feature fusion, the Gaussian-Geo module leverages Gaussian kernel rendering on sparse LiDAR depth maps to enhance 2D image features with dense geometric priors, and the Semantic-Aware module enriches LiDAR voxels with semantic context via deformable cross-attention; (2) In the late-stage voxel fusion, the Adaptive Fusion (AF) module dynamically balances voxel features across modalities, while the High Classification Confidence Voxel Fusion (HCCVF) module resolves semantic inconsistencies using self-attention-based refinement. Experiments on two large-scale benchmarks demonstrate state-of-the-art performance. On nuScenes-OpenOccupancy, MS-Occ achieves an Intersection over Union (IoU) of $32.1\%$ and a mean IoU (mIoU) of $25.3\%$, surpassing the state-of-the-art by $0.7\%$ IoU and $2.4\%$ mIoU. Furthermore, on the SemanticKITTI benchmark, our method achieves a new state-of-the-art mIoU of $24.08\%$, robustly validating its generalization capabilities. Ablation studies further confirm the effectiveness of each individual module, highlighting substantial improvements in the perception of small objects and reinforcing the practical value of MS-Occ for safety-critical autonomous driving scenarios.
@@ I Introduction
3D semantic occupancy is vital for autonomous driving, enabling a comprehensive understanding of complex environments for downstream tasks such as vehicle planning and control. Unlike traditional 3D object detection, occupancy perception captures irregular and non-rigid objects beyond predefined categories. Vision-based 3D semantic occupancy perception has shown notable performance across various technical paradigms, including monocular setups [monoscene], surround-view camera systems [surroundocc, bevformer, FB-BEV], and efficient architectural designs [fullySparseOcc, FastOcc]. However, vision-based methods face inherent challenges like inaccurate depth estimation, limited geometric fidelity, and sensitivity to lighting conditions, undermining robustness in real-world driving environments. These challenges highlight the importance of integrating LiDAR and camera data to improve occupancy perception. Although 4D radar has been explored in recent studies [OmniHDScenes, Doracamom] as a supplement to mitigate some of these issues, it remains constrained by limitations such as point cloud noise and sparsity.
@@
Figure 1: Comparison of occupancy fusion stages in encoder design. (a) Middle-stage fusion (e.g., [occloff]), where 2D image features are directly fused with 3D LiDAR voxel features. (b) Late-stage fusion, including methods such as [wang2023openoccupancy, Co-Occ, OccGen, EFFOcc], which explicitly lift 2D image features to 3D voxel features (e.g., using [lss]), followed by 3D fusion in voxel space. (c) Multi-stage fusion, as proposed in our MS-Occ, integrates both middle and late-stage fusion within a hybrid architecture.
@@
LiDAR-camera integration is a key research focus in multi-modal sensor fusion. Multi-modal 3D semantic occupancy perception can be classified into four fusion stages based on the timing of feature integration [MultiModalObjectDetection, multilayer_depthEstimationRC]: early-stage fusion, middle-stage fusion, late-stage fusion, and multi-stage fusion. Early-stage fusion involves the combination of raw or preprocessed sensor data, allowing for full utilization of the original information. However, it is prone to misalignment caused by modality-specific limitations and calibration errors. As a result, most existing methods prioritize middle or late-stage fusion, as illustrated in Fig. 1(a) and (b).
@@
Middle-stage fusion integrates 2D image features with 3D LiDAR features through projection-based fusion $E_{\text{fusion}}^{\text{3D}\rightarrow\text{2D}}$ or cross-attention mechanisms $E_{\text{fusion}}^{\text{2D}\rightarrow\text{3D}}$. For instance, OccLoff [occloff] employs cross-attention to fuse 2D image features with LiDAR voxel features. Nonetheless, middle-stage fusion methods often fail to fully exploit the strengths of each modality within the 3D voxel space. In contrast, late-stage fusion combines 3D image voxel features, explicitly lifted from 2D image features ($E_{\text{imgs}}^{\text{3DLift}}$), with 3D LiDAR voxel features in voxel space ($E_{\text{fusion}}^{\text{3D}}$). Representative examples include M-CONet [wang2023openoccupancy] and Co-Occ [Co-Occ], both of which enhance performance through late-stage fusion. OpenOccupancy [wang2023openoccupancy] presents a novel LiDAR-camera fusion baseline that leverages the complementary advantages of the sensors and establishes the first benchmark for fusion-based surrounding occupancy perception. Co-Occ [Co-Occ] further refines LiDAR features by incorporating neighboring camera features using a KNN search strategy in GSFusion.
@@
Despite their success, these approaches perform fusion solely in voxel space, overlooking the hierarchical nature of 2D image features. This omission introduces geometric and semantic biases during voxel grid construction, thereby limiting the alignment of multi-modal features in the 3D voxel domain. To address these limitations, multi-stage fusion, illustrated in Fig. 1(c), combines both middle- and late-stage fusion strategies. This method leverages both 2D visual and 3D LiDAR features for a more robust alignment of semantic richness and geometric precision. Nevertheless, multi-stage fusion remains underexplored in existing research.
@@
This study introduces MS-Occ, to our knowledge, the first LiDAR-camera fusion framework for 3D semantic occupancy prediction using multi-stage fusion. Initially, features are extracted from both LiDAR point clouds and surround-view images. These features are then enhanced through middle-stage feature fusion, where 2D camera features are enriched with geometric information using the proposed Gaussian-Geo module, and LiDAR features are augmented with semantic cues via the Semantic-Aware module. In the Gaussian-Geo module, a Gaussian kernel is applied to produce a denser LiDAR heatmap, which interacts with the original 2D image features to generate geometrically enhanced camera features. These enhanced features are subsequently fused with 3D LiDAR features using deformable attention in the semantic-aware module, thereby injecting semantic context into the LiDAR representations.
@@
Following this middle-stage fusion, a late-stage voxel fusion is performed to further refine the fused voxel features from both modalities. This stage comprises two complementary components: the High Classification Confidence Voxel Fusion (HCCVF) module and the Adaptive Fusion (AF) module. The AF module initially performs adaptive fusion of voxel features, following the fusion method of M-CONet [wang2023openoccupancy], to integrate voxel representations from both LiDAR and camera streams. The innovative HCCVF Module then selectively refines high-confidence features from each modality through an attention mechanism, thereby mitigating semantic ambiguities between them. The final fused voxel features are passed through a prediction head to generate the 3D semantic occupancy output.
@@
  • A novel LiDAR-camera multi-stage fusion framework for 3D semantic occupancy prediction is proposed. To our knowledge, this is the first multi-stage fusion architecture for this task. The framework efficiently processes LiDAR and camera data to generate 3D voxel representations that compensate for modality-specific limitations.
  • Four dedicated modules are introduced for unified and fine-grained perception. The Gaussian-Geo module enriches camera features with dense geometric priors. The Semantic-Aware module transfers image semantics to LiDAR voxels through deformable attention. In the late-stage, the HCCVF and AF modules are integrated: HCCVF resolves semantic ambiguities via self-attention, while the AF module dynamically balances geometric and semantic voxel features.
  • Extensive experiments on the nuScenes-OpenOccupancy and SemanticKITTI benchmarks demonstrate that MS-Occ outperforms existing state-of-the-art methods, achieving superior geometric and semantic accuracy, with significant improvements in safety-critical categories.
@@ II-A Vision-based 3D Occupancy Prediction
Vision-based occupancy perception has seen significant recent advances. Transforming 2D image features into 3D representations (e.g., Bird’s-Eye View (BEV) features [yu2023flashocc], Tri-Perspective View (TPV) features [TPVFormer], and voxel features [surroundocc]) is critical for this task. Three dominant approaches have emerged for this transformation [occ_survey]: (1) The projection-based method [monoscene, miao2023occdepth] employs perspective projection models to establish geometric correspondences by projecting voxel centroids from 3D space onto 2D image planes. (2) The depth prediction method [zhang2023occformer, yu2023flashocc, FastOcc], derived from the Lift-Splat-Shoot (LSS) paradigm, uses estimated depth distributions to lift 2D features into 3D space. (3) The cross-attention method [surroundocc, wang2024panoocc, TPVFormer] facilitates feature interaction between 3D voxel grids and 2D feature maps through deformable attention [DeformableDETR]. Recent works such as SparseOcc [fullySparseOcc] and EFFOcc [EFFOcc] have also improved computational efficiency. However, vision-based 3D occupancy perception faces fundamental limitations due to sensitivity to illumination changes, inherent depth estimation errors, and compromises in geometric precision.
@@ II-B LiDAR-Camera Fusion for 3D Occupancy Prediction
LiDAR-camera fusion has shown exceptional performance via four primary strategies [occ_survey]: (1) Concatenation-based methods [occfusion, Co-Occ], which concatenate 3D feature volumes from distinct modalities. (2) Summation-based approaches [wang2023openoccupancy, OccGen], which perform element-wise additive operations to combine geometrically aligned sensor representations. (3) Attention-based methods [occloff], inspired by SparseFusion [sparseFusion] and TransFusion [TransFusion], which apply cross-modality attention mechanisms. (4) Projection-based methods [mseg3d_cvpr2023, Huang2020EPNet], which align LiDAR points with image pixels via calibrated projection. Recent research fuses RGB semantics with depth for dense scene reconstruction [rosinol2021kimera, ilabel, zheng2024semantic], further highlighting the shift toward semantically enriched geometry.
However, these strategies primarily focus on single middle- or late-stage fusion mechanisms, neglecting fusion timing and thereby constraining dynamic alignment of representations across modalities. To address these challenges, we introduce a two-stage enhancement and fusion strategy. In the middle stage, camera features are enriched with geometric priors through lightweight Gaussian kernel rendering (Gaussian-Geo module), which provides dense cues without requiring heavy separate depth networks, while LiDAR features are augmented with semantic context using efficient deformable attention (Semantic-Aware module), which adaptively queries salient information more efficiently than global attention. In the late stage, voxel-level fusion is performed by first applying adaptively weighting (AF module) and subsequently refining the result through targeted self-attention (HCCVF module) refines high-confidence or conflicting voxels to resolve semantic inconsistencies, a more robust alternative than conventional summation or simple concatenation that treats all voxels equally. The proposed MS-Occ leverages modality-specific features at multiple fusion stages to facilitate cross-modal alignment and enhance the quality of fused representations.
@@ III Method
The overall architecture is illustrated in Fig. 2. Modality-specific features are acquired using 2D and 3D encoders. The encoded 2D image features, are denoted as $F_{\mathrm{I}}^{(m,i)}\in\mathbb{R}^{H_{\mathrm{I}}\times W_{\mathrm{I}}\times C_{\mathrm{I}}}$ , where $m$ and $i$ indicate the $m$-th camera view and $i$-th image scale, respectively. Here, $H_{\mathrm{I}}$ and $W_{\mathrm{I}}$ represent the spatial dimensions of the image feature map, and $C_{\mathrm{I}}$ denotes the feature channel dimension. The sparse 3D LiDAR voxel features are denoted as $F_{\mathrm{L}}\in\mathbb{R}^{D\times H\times W\times C}$, where $D$, $H$, $W$ are the depth, height, and width of the voxel grid, and $C$ is the number of feature channels for each voxel.
@@ III-A1 Gaussian-Geo Module
The Gaussian-Geo module is designed to incorporate geometric priors through Gaussian kernel rendering, drawing inspiration from recent approaches [gaussianocc, gaussianFormer] that utilize 3D Gaussians [3DGS] for scene representation, as illustrated in Fig. 3. Specifically, depth maps are first obtained from LiDAR point clouds. A straightforward approach involves directly projecting point clouds onto images using intrinsic and extrinsic calibration parameters to generate depth maps. However, due to the inherent sparsity of LiDAR data, the resulting depth maps are typically incomplete and lack detailed features. To mitigate this limitation, a Gaussian kernel rendering technique is introduced to enhance the projected depth maps across multi-view images, resulting in more informative depth representations $Y\in[0,U]^{H_{\mathrm{C}}\times W_{\mathrm{C}}\times 1}$, where $U$ denotes the LiDAR point depth value, $H_{\mathrm{C}}$ and $W_{\mathrm{C}}$ represent the height and width of the original camera image, respectively. The raw LiDAR points are splatted onto the image planes of multiple views. For each hit position $p_{g}\in\mathbb{R}^{\mathrm{2}}$ as a center, a Gaussian kernel with the corresponding depth value is generated in the image plane. The final depth value at position $q\in\mathbb{R}^{\mathrm{2}}$ is computed as:
@@
$Y_{q}=\underset{g\in\mathcal{G}_{\text{hit}}}{\min}\left(U^{g}\mathtt{exp}\left(-\frac{\left(q-p_{g}\right)^{2}}{\mathrm{2\sigma^{2}}}\right)\right),$ (1)
@@
where $\sigma$ controls the spread of the kernel, $\mathcal{G}_{\text{hit}}$ represents the set of overlapping Gaussian kernels at position $q$, and $g$ indexes the overlapping kernels.
Figure 3: Illustration of intermediate outputs from the proposed Gaussian-Geo module. The LiDAR point cloud is densified using Gaussian kernel, and the resulting geometric information is subsequently transferred to the image to enhance its spatial representation.
@@
In regions with multiple overlapping kernels, the minimum value is selected to generate a denser and more continuous multi-view depth map. The enhanced depth maps $Y$ are then passed through a shared encoder $\varphi\left(\cdot\right)$ to extract depth features $F_{\mathrm{E}}^{(m,i)}\in\mathbb{R}^{H_{\mathrm{I}}\times W_{\mathrm{I}}\times\mathcal{A}}$ , where $\mathcal{A}$ denotes the dimension of the depth features. These features are concatenated with the corresponding 2D image features $F_{\mathrm{I}}^{(m,i)}$, and the composite features are processed by a depth fusion network $\psi\left(\cdot\right)$ to obtain depth-aware 2D image features $F_{\mathrm{E2I}}^{(m,i)}\in\mathbb{R}^{H_{\mathrm{I}}\times W_{\mathrm{I}}\times C_{\mathrm{I}}}$, thereby enhancing the geometric understanding of the image content. The overall process can be formulated as:
$F_{\mathrm{E2I}}^{(m,i)}$ $=\psi\left(\mathtt{Concat}\left[F_{\mathrm{E}}^{(m,i)},F_{\mathrm{I}}^{(m,i)}\right]\right),$ (2)
$F_{\mathrm{E}}^{(m,i)}$ $=\varphi\left(Y\right).$ (3)
@@
Ultimately, the depth-aware 2D image features $F_{\mathrm{E2I}}^{(m,i)}$ are fed into the LSS [lss] module, which transforms them into 3D voxels $V_{\mathrm{I}}\in\mathbb{R}^{D\times H\times W\times C}$ from image modality, matching the dimensions of the LiDAR voxel features. This transformation enriches geometric-aware 2D image features, yielding a more comprehensive and spatially accurate scene representation.